OpenStudy (anonymous):

Prove that this equation is an identity: {cosine(alpha)-sine(alpha)}^2=1-2 sine(alpha)cosine(alpha)

OpenStudy (anonymous):

\[(cosx - sina)^2 = cos^2x + sin^2x - 2*sinx*cosx\]

OpenStudy (anonymous):

you DIDN'T ANSWER THE QUESTION!!!!!

OpenStudy (anonymous):

I did !!!

OpenStudy (anonymous):

no.... you just restated it and added more....

OpenStudy (anonymous):

\[sin^2x + cos^2x = 1 \]

OpenStudy (anonymous):

\[\text{^ ^ Now try to do on your own}\]

OpenStudy (anonymous):

i did... i kinda got it not really

myininaya (myininaya):

\[(\cos(x)-\sin(x))^2=\cos^2(x)+\sin^2(x)-2\cos(x)\sin(x)=1-2\cos(x)\sin(x)\] but ishaan already stated this